Eggz, Inc., is considering the purchase of new equipment that will allow the company to collect loose hen feathers for sale. The equipment will cost $425,000 and will be eligible for 100 percent bonus depreciation. The equipment can be sold for $25,000 at the end of the project in 5 years. Sales would be $275,000 per year, with annual fixed costs of $47,000 and variable costs equal to 35 percent of sales. The project would require an investment of $25,000 in NWC that would be returned at the end of the project. The tax rate is 22 percent and the required return is 9 percent.


Calculate the NPV of this project.
1
Expert's answer
2020-10-21T10:17:09-0400

The NPV of this project is:

"NPV = -(425,000 + 25,000) + \frac{275*(1 - 0.35)*(1 - 0.22) + 80,000 - 47,000}{1 + 0.09} + \frac{275*(1 - 0.35)*(1 - 0.22) + 80,000 - 47,000}{(1 + 0.09)^2} + \frac{275*(1 - 0.35)*(1 - 0.22) + 80,000 - 47,000}{(1 + 0.09)^3} + \frac{275*(1 - 0.35)*(1 - 0.22) + 80,000 - 47,000}{(1 + 0.09)^4} + \frac{275*(1 - 0.35)*(1 - 0.22) + 80,000 - 47,000 + 25,000 + 25,000}{(1 + 0.09)^5} = -288,602.62."